Find the maximum value of


x^2y^3z^4


subject to the condition


x+y+z=5

1
Expert's answer
2022-01-17T16:06:03-0500

"\text{Let, } a=x^2y^3z^4
\Rightarrow \text{ln } a=2\text{ ln } x+3\text{ ln } y+4\text{ ln } z\\"

Let

"P=ln (a)+\lambda (x+y+z-5)\\
\therefore \frac{\partial P}{\partial x}=\frac{2}{x}+\lambda,~
\frac{\partial P}{\partial y}=\frac{3}{y}+\lambda, ~
\frac{\partial P}{\partial z}=\frac{4}{z}+\lambda"

If we maximizing a implies maximizing P.

i.e. "\frac{\partial P}{\partial x}=\frac{\partial P}{\partial y}=\frac{\partial P}{\partial z}=0"

"\Rightarrow \lambda=-\frac{2}{x}=-\frac{3}{y}=-\frac{4}{z}\\
\therefore \frac{2}{x}=\frac{3}{y}=\frac{4}{z}= \frac{2+3+4}{x+y+z}=\frac{9}{5} \\
\texttt{[As we are doing this under the constraint x+y+z=5]}\\"

"\Rightarrow x=\frac{10}{9}, y=\frac{15}{9}, z=\frac{20}{9}\\
\text{Hence the maximum value of $a$ is }\\
\max_{x,y,z}a=(\frac{10}{9})^2\times(\frac{5}{3})^3\times(\frac{10}{9})^4=\dfrac{2\times10^9}{3^{15}}"